o Probability Rules 10. The Committee on Student Life did a survey of 417 students regarding satisfaction with Student Government and class standing. The results follow: Freashman Sophomore Junion Total Not Satisfied Neutral Satisfied23 Total 17 61 32 3816 A3 65 80 0311A98 16 17 101 Assume the sample represents the entire population o students. Find the probability that a student selected at random is: (a) Satisfied (b) Satisfied, given the student is a Senior (c) Neutral (d) Neutral and freshman (e) Neutral, given the student is a Freshmarn (f) Senior, given satisfied (g) Are the events, Freshman and neutral independent or not? Explain SAMSUNG
831 0930 0339 03 Probability Rules 10. The Committee on Student Life did a survey of 417 students regarding satisfaction with Student Government and class standing. The results follow Freashman Sophomore Junior Senior Total Not Satisfied17 Neutral 61 Satis Total 19 35 19 23 12 71 32 38 166 43 65 1S0 23 101 103 8115 417 Assume the sample represents the entire population of students. Find the probability that a student selected at random is: (a) Satisfied (b) Satisfied, given the student is a Senior (e) Neutral (d) Neutral and freshman (e) Neutral, given the student is a Freshman (f) Senior, given satistied (g) Are the events, Freshman and neutral independent or not? Explain SAMSUNG
Answer #1

A) P(satisfied) = 180/417 = 0.4317

B) P(satisfied | senior) = P(satisfied and senior)/P(senior)

= 65/115 = 0.5652

C) P(neutral) = 166/417 = 0.3981

D) P(neutral and freshman) = 61/417 = 0.1463

E) P(neutral | freshmen) = P(neutral and freshman)/P(fresh men) = 61/101 = 0.6040

F) P(senior | satisfied) = P(senior and satisfied)/P(satisfied)

= 65/180 = 0.3611

G) P(freshmen and neutral) = 61/417 = 0.1463

P(freshmen) * P(neutral) = 101/417 * 166/417 = 0.0964

Since P(freshmen and senior)  \neq P(freshmen) * P(senior), so these two events are not independent.